# Magnetic hardness of hexagonal and orthorhombic Fe3C, Co3C, (Fe-Co)3C, and their alloys with boron, nitrogen, and transition metals: A first-principles study

### v2

@@ I Introduction
When discussing Fe3C and Co3C alloys for use as permanent magnets, we will focus primarily on two intrinsic properties of the materials, magnetocrystalline anisotropy energy (MAE) and Curie temperature (). The experimental value of the uniaxial magnetocrystalline anisotropy constant of ortho-Fe3C is a moderate 0.405 MJ m-3 at 5 K [20]. Whereas, the measured Curie temperature for ortho-Fe3C is 481 K [21]. The state of the knowledge on the properties of ortho-Fe3C (cementite) is summarized by Bhadeshia [22]. The ortho-Co3C shows an experimental MAE of 0.740.1 MJ m-3 and Curie temperature of 650 K [23]. Additionally, Co3C carbide nanocrystals have been confirmed to possess hard magnetic properties [12]. Enhanced magnetocrystalline anisotropy has been also observed in Co-C nanowires, Co-C nanoparticles, and polyphase Fe-Co-C polycrystalline samples [24, 23, 15]. The melt-spun Fe2CoC alloys show elevated magnetocrystalline anisotropy constant () of 0.91 MJ m-3 and magnetic moment of 1.23 atom-1 (at 10 K) [17]. In contrast, an exceptionally high of 4.6 MJ m-3 has been measured for Fe2CoC nanoparticles [15].
@@ II Calculations details
where , , , and are the total energies of the supercell and the crystals of iron, carbon (graphite), and transition metal (X), respectively. To determine the total energies, we prepared and optimized the geometry of 3, 4, and 5 crystals. The formation energies of hexa-Co11X1C4 compositions were calculated analogicaly.
@@
where is total spin magnetic moment, is total orbital magnetic moment, and is unit cell volume. It is assumed that defines the range of semi-hard magnetic materials, while denotes hard magnetic materials.
Curie temperature () of (Fe0.916X0.084)3C alloys with the 3, 4, and 5 transition metals (X) were calculated using mean-field theory [66, 67]. The (Fe0.916X0.084)3C concentration, modeled here with coherent potential approximation (CPA), was chosen to match the Fe11X1C4 concentration, modeled with supercell method. We used the:
equation. Here, and are the total energies of the paramagnetic and ferromagnetic solutions, is Boltzmann’s constant, and is the number of magnetic atoms in formula. For Fe3C we assume equal 3. To model the paramagnetic configuration, we used the disordered local moment (DLM) approach [68] with the coherent potential approximation (CPA) [69]. Scalar-relativistic DLM-CPA calculations were conducted using the FPLO5 code, which is the most recent version of the FPLO code that incorporates CPA. However, the FPLO5 code lacks the PBE-GGA implementation, forcing the use of the LDA-PW92 exchange-correlation potential for CPA calculations [63].
| Vxc | hexa-Fe3C | hexa-Co3C | ortho-Fe3C | ortho-Co3C | ||||
| MAE | MAE | MAE | MAE | |||||
| PBE [17] | 5.92 | 0.57 | 2.76 | 0.02 | 5.52 | 0.05 | 3.12 | 0.81 |
| PBE | 6.05 | 0.68 | 2.95 | -0.15 | 5.56 | 0.11 | 3.16 | 0.67 |
| vBH | 5.40 | 0.55 | 2.48 | 0.26 | 4.88 | 0.24 | 2.76 | 0.81 |
| PZ | 5.40 | 0.55 | 2.48 | 0.26 | 4.88 | 0.24 | 2.77 | 0.81 |
| PW92 | 5.66 | 0.71 | 2.67 | 0.02 | 5.16 | 0.35 | 2.89 | 0.88 |
| x-only | 6.36 | 0.41 | 3.20 | 0.03 | 5.88 | 0.05 | 3.38 | 0.72 |
@@ III.1 Fe3C and Co3C compounds
The calculated formation energies are positive for all four compounds considered in this section (hexa-Fe3C, hexa-Co3C, ortho-Fe3C, and ortho-Co3C), indicating their instability, which agrees with experimental formation enthalpies [70] and previous computational results [17, 33, 19]. Orthorhombic phases have slightly lower energies than hexagonal ones, suggesting that Fe3C and Co3C will prefer the cementite-type structure. The formation energies are also relatively low, giving the possibility of stabilizing the compounds with additional factors such as alloying or the presence of a second phase. Nanoparticles of ortho-Co3C were successfully synthesized [13, 71] and the bulk form of ortho-Co3C has been obtained under high pressure (above 4.8 GPa) [19]. Another method of stabilizing ortho-Co3C is its alloying with a stable isostructural Co3B phase [72].
Figure 2 presents the densities of states (DOS) calculated for orthorhombic Fe3C and Co3C. Densities of states for hexagonal phases look very similar to orthorhombic, and we decide to not show them. The valence band of ortho-Fe3C is dominated by Fe 3 and C 2 orbitals, between which a clear hybridization is visible. The spin polarization of the valence band is noticeably stronger for Fe3C.
@@
In summary, from the perspective of magnetic hardness, the ortho-Co3C phase deserves the most attention. Calculations suggest that magnetic hardness close to unity should persist not only at 0 K, but also at room temperature when the magnetic moment decreases. Having determined the magnetic hardness of Fe3C and Co3C compounds, in the next section we will address the full ranges of intermediate concentrations of (Fe,Co)3C alloys.
@@
Since Fe and Co have consecutive atomic numbers of 26 and 27, their alloys can be modeled using a virtual atom with a fractional atomic number determined by the concentration of the elements. Our goal is to study how the MAE of the system changes with Co concentration and identify promising compositions with elevated MAE. In our attempt we are motivated by previous studies, such as for FeCo, (Fe,Co)2B and (Fe,Co)16C alloys [77, 78, 79], which showed higher MAE values for intermediate compositions than for compounds at the extremes of the Fe/Co range. Moreover, we will determine how the spin and orbital magnetic moments change with Co concentration.
@@
At first, we analyze the concentration dependence of MAE of the hexa-(Fe,Co)3C phase, see Fig. 4(a) and Table 4. We observe both positive and negative MAE values, indicating uniaxial and in-plane magnetocrystalline anisotropies, respectively. The highest MAE of 1.07 MJ m-3 was obtained for a concentration of 20% Co, and the minimum MAE of -1.46 MJ m-3 for 70% Co. This trend is not surprising, as an analogous course of MAE() with a positive maximum and a negative minimum was observed for the mentioned before (Fe1-xCox)2B and (Fe1-xCox)16C alloys [78, 79].
@@
For ortho-(Fe,Co)3C alloys, the Co concentration range is divided into regions: from 0% to 35%, from 35% to 75%, and from 75% to 100%, with the directions of magnetic easy axes [010], [001], and [100], respectively, see Fig. 4(b). The highest MAE value of 0.82 MJ m-3 was obtained for a concentration of 90% Co. Additionally, for orthorhombic alloys, we show the energy difference between two higher energies (DE) complementing the MAE. The low DE32 values, as observed for Co-rich phases, strengthen the uniaxial character of the magnetocrystalline anisotropy of these systems.
Wu et al. synthesized a sample of Fe2CoC and measured the MAE for it at 10 K as 0.96 MJ m-3 [17]. In contrast, our calculations for about 1/3 Co concentration indicate MAE close to zero. Such a discrepancy may be due to the limitations of the VCA method used and the high value of DE32, but also to the presence of other phases in the experimental sample and the influence of the microstructure on the result measured. Nevertheless, the relatively high experimental MAE value encourages further exploration of magnetically hard materials in the cementite family.
@@
The total spin magnetic moment () is highest for the terminal Fe3C phases for both hexagonal and orthorhombic structures, and decrease with Co concentration, see Figs. 4(c,d). We do not observe a pronounced maximum in the intermediate range. This behavior is similar to that observed, for example, for Fe/Co monoborides [64].
@@
Hexagonal (Fe,Co)3C alloys with Co concentration from 0 to 30% are magnetically semi-hard (0.5 < < 1), see Fig.4(g). For negative values of MAE, seen in the center of the concentration range, we assumed = 0, see Table 4. The hexagonal carbide with 90% Co indicates = 1.13, which defines it as magnetically hard.
For orthorhombic alloys, see Fig.4(h), semi-hard magnetic properties are observed for most of Co concentrations above 50%.
@@ III.3.1 Formation energy
The calculated formation energy of hexa-Fe3C is about 0.085 eV atom-1 and is slightly higher than that for ortho-Fe3C (0.055 eV atom-1). The situation is similar when comparing trends for hexa-Fe3C and ortho-Fe3C alloys with transition metals. The calculated formation energy of ortho-Fe3C (0.055 eV atom-1) stays in a good agreement with previous theoretical and experimental results (0.05–0.08 eV atom-1) [33, 45, 50]. The plots of dependencies of formation energies on 3 and 4 transition metal substitutions in ortho-Fe3C are almost identical to the previously mentioned theoretical results [33, 50]. Although the formation energies of the alloys studied provide valuable information, the idealized single-phase nature of the models means that the actual impact of additives is best verified experimentally. Bulk ortho-Fe3C has been tested by Umemoto et al. [27], who found that additions of Cr, Mn, V and Mo additions stabilize the cementite, whereas Ti, Ni, and Si destabilize it. Furthermore, it has been determined that the substitution of 5 at.% Ti leads to the appearance of additional phases of TiC and -Fe [27]. The formation energies of ortho-Co3C alloys are higher than those of ortho-Fe3C, however, still relatively low.
@@ III.3.2 Curie temperature
From the standpoint of potential applications of the material in permanent magnets, it is important that its Curie temperature far exceeds the room temperature. For metastable hexa-Fe3C, the Curie temperature is difficult to determine and only an approximate experimental value of K is given [81]. Whereas, off-stoichiometric hexa-Fe2.2C has of about 723 K [82]. Experimental Curie temperatures of ortho-Fe3C and ortho-Co3C are about 481 K [21] and 650 K [23], respectively. In Fig. 6 we present the mean-field theory Curie temperatures () calculated for hexa-(Fe0.916X0.084)3C, ortho-(Fe0.916X0.084)3C, and ortho-(Co0.916X0.084)3C alloys with transition metals X. The transition metal concentration of 0.084 was chosen to reproduce the concentration of transition metal element X in Fe11X1C4 models.
The calculated for hexa-Fe3C is 1040 K, which is much higher than the mentioned experimental estimation ( K). The observed difference in may come from both the known overestimation of in the mean-field theory and the underestimation of the experimental value. In Fig. 6 we see, that the dependencies of on the atomic number of substitution have a clear sinusoidal trend and that each substitution reduces the of the parental hexa-Fe3C compound. Nevertheless, all the considered hexa-Fe3C + X alloys show relatively high .
Calculated of ortho-Fe3C is equal to 516 K, in good agreement with experimental value of 481 K [21]. Calculated trends for transition metal elements suggest that all substitutions, except Co, will significantly reduce the of the alloy. Whereas, the elements 4 and 5 will have stronger effect than the elements 3. The transition metal elements on the left side of the periods reduce to zero, which means the absence of magnetic ordering. Previous experiments confirm the reduction of Curie temperature after alloying of cementite with Cr, Mn, or Mo [25, 26]. However, it has been measured that up to about 2% Ni raises the Curie temperature of the cementite [25]. The effect of lowering can make it difficult or impossible to use most transition metals as additives to affect MAE and magnetic hardness. The solution could be, for example, a combination of transition metal alloying on the Fe site with B alloying on the C side. Indeed, the experiment showed a linear growth in up to 851 K with increasing B concentration in ortho-Fe3(B,C) borocarbides [21].
Calculated of ortho-Co3C is equal to 458 K, whereas experimental of Co3C-based nanoparticles have been determined as 510 K [83], 563 K [84], or 650 K [23]. Calculations indicate that only Fe substitution raises the of the ortho-Co3C-based system, while all other elements considered lower it. The determined values below and well below 400 K suggest the impossibility of using most ortho-Co3C + X alloys as permanent magnets. The solution, as we suggested above, could be the alloying of B at the C position, since, as shown experimentally, the of ortho-Co3B is much higher than the value for ortho-Co3C and is 750 K [16].

### v3

@@ I Introduction
When discussing Fe3C and Co3C alloys for use as permanent magnets, we will focus primarily on two intrinsic properties of the materials, magnetocrystalline anisotropy energy (MAE) and Curie temperature (). The experimental value of the uniaxial magnetocrystalline anisotropy constant of ortho-Fe3C is a moderate 0.405 MJ m-3 at 5 K [20]. Whereas, the measured Curie temperature for ortho-Fe3C has been measured as 481 K [21] and 483 K [22]. The state of the knowledge on the properties of ortho-Fe3C (cementite) is summarized by Bhadeshia [23]. For ortho-Co3C, the measured Curie temperatures are 498 K [24], 510 K [25], 563 K [26], and 650 K [27]. The results were determined for nanoparticles, and their large dispersion indicates a significant effect of structurization on the magnetic properties of the system. The ortho-Co3C shows an experimental MAE of 0.740.1 MJ m-3 [27]. Additionally, Co3C carbide nanocrystals have been confirmed to possess hard magnetic properties [12]. Enhanced magnetocrystalline anisotropy has also been observed in Co-C nanowires, Co-C nanoparticles, and polyphase Fe-Co-C polycrystalline samples [28, 27, 15]. The melt-spun Fe2CoC alloys show elevated magnetocrystalline anisotropy constant () of 0.91 MJ m-3 and magnetic moment of 1.23 atom-1 (at 10 K) [17]. In contrast, an exceptionally high of 4.6 MJ m-3 has been measured for Fe2CoC nanoparticles [15].
@@ II Calculations details
where , , , and are the total energies of the supercell and the crystals of iron, carbon (graphite), and transition metal (X), respectively. To determine the total energies, we prepared and optimized the geometry of 3, 4, and 5 crystals. The formation energies of hexa-Co11X1C4 compositions were calculated analogically. The formation energies are determined for the ground state, that is, at 0 K.
@@
where is total spin magnetic moment, is total orbital magnetic moment, and is unit cell volume. While denotes hard magnetic materials, it is assumed that defines the range of semi-hard magnetic materials [11]. To use the semi-hard materials as permanent magnets, the condition that should be met [11].
Curie temperature () of (Fe0.916X0.084)3C alloys with the 3, 4, and 5 transition metals (X) were calculated using mean-field theory [70, 71]. Here, mean-field theory is an approximation for fluctuations in configuration of magnetic moment orientations, where all interactions with any body are replaced by an averaged interaction [70]. When calculating the mean-field Curie temperature (), a coherent potential approximation (CPA) procedure can be used to average and iterate to self-consistency [72, 70]. The (Fe0.916X0.084)3C concentration, modeled here with CPA, was chosen to match the Fe11X1C4 concentration, modeled with supercell method. We used the:
equation. Here, and are the total energies of the paramagnetic and ferromagnetic solutions, is Boltzmann’s constant, and is the number of magnetic atoms in formula. For Fe3C we assume equal 3. To model the paramagnetic configuration, we used the disordered local moment (DLM) approach [73] with the coherent potential approximation (CPA) [72]. In the DLM approach, the local magnetic moments can rotate, and the self-consistent calculations are conducted for the given arrangement of them. Numerous magnetic moment orientations are calculated and averaged. The paramagnetic state is determined by taking the total magnetic moment equal to zero. A ferromagnetic (FM) ground state, on the other hand, is defined as an equilibrium configuration of parallel collinear magnetic moments. Scalar-relativistic DLM-CPA calculations were conducted using the FPLO5 code, which is the most recent version of the FPLO code that incorporates CPA. However, the FPLO5 code lacks the PBE-GGA implementation, forcing the use of the LDA-PW92 exchange-correlation potential for CPA calculations [67].
@@ III.1 Fe3C and Co3C compounds
The experimental standard enthalpy of formation () for Fe3C is kJ mol-1 ( eV atom-1) and for Co3C is kJ mol-1 ( eV atom-1) [74]. Positive enthalpy values indicate instability. Similarly, the previously calculated corresponding formation energies are positive and are 0.016 eV atom-1 [54] and 0.07 eV atom-1 [49] for ortho-Fe3C and 0.05 eV atom-1 for ortho-Co3C [19]. The formation energies calculated in this work are positive for all four compounds considered in this section and equal to 0.085 eV atom-1 for hexa-Fe3C, 0.70 eV atom-1 for hexa-Co3C, 0.055 eV atom-1 for ortho-Fe3C, and 0.14 eV atom-1 for ortho-Co3C. Orthorhombic phases have lower energies than hexagonal ones, suggesting that Fe3C and Co3C will prefer the cementite-type structure. The formation energies are also relatively low, giving the possibility of stabilizing the compounds with additional factors such as alloying or the presence of a second phase. Nanoparticles of ortho-Co3C were successfully synthesized [13, 75] and the bulk form of ortho-Co3C has been obtained under high pressure (above 4.8 GPa) [19]. Another method of stabilizing ortho-Co3C is its alloying with a stable isostructural Co3B phase [76].
For the application of the material in permanent magnets, it is important that its Curie temperature far exceeds the room temperature. For metastable hexa-Fe3C, the Curie temperature is difficult to determine and only an approximate experimental value of K is given [77]. Whereas, off-stoichiometric hexa-Fe2.2C has of about 723 K [78]. The calculated for hexa-Fe3C is 1040 K, which is much higher than the mentioned experimental estimation ( K). The observed difference in may come from both the known overestimation of in the mean-field theory and the underestimation of the experimental value. Because, as our calculations indicate, the hexa-Co3C phase has a much higher formation energy than the other phases under consideration, it is much more difficult to stabilize, and we have not found literature experimental results on both its synthesis and Curie temperature. The value we determined for hexa-Co3C is 310 K.
The experimental Curie temperatures of ortho-Fe3C is equal to about 481 K [21] or 483 K [22], while our calculations indicate 570 K, overestimating the experimental value by about 20%, which is the typical behavior of the mean-field results [79, 80, 81]. The experimental Curie temperature for ortho-Co3C can be estimated from the results for ortho-Co3C-based nanoparticles, but these are ambiguous and depend on the specific properties of the nanoparticles. For ortho-Co3C-based nanoparticles, we encounter the following Curie temperatures: 498 K [24], 510 K [25], 563 K [26], and 650 K [27]. calculated by us for ortho-Co3C is equal to 458 K. The observed discrepancy between the mean-field result and measurements confirms the significant effect of nanostructuring on the magnetic properties of ortho-Co3C nanoparticles.
Figure 2 presents the densities of states (DOS) calculated for orthorhombic Fe3C and Co3C. The valence band of ortho-Fe3C is dominated by Fe 3 and C 2 orbitals, between which a clear hybridization is visible. The spin polarization of the valence band is noticeably stronger for Fe3C. For densities of states for hexagonal phases, see Supplementary Material.
@@
In summary, from the perspective of magnetic hardness, the ortho-Co3C phase deserves the most attention. Calculations suggest that magnetic hardness close to unity should persist not only at 0 K, but also at room temperature when the magnetic moment decreases. In addition, previous experiments show high values of the magnetic anisotropy constant for ortho-Co3C nanoparticles and Curie temperatures above 500 K. However, the relatively high price of cobalt works against the potential applications of this system. Having determined the magnetic hardness of Fe3C and Co3C compounds, in the next section we will address the full ranges of intermediate concentrations of (Fe,Co)3C alloys.
@@
Since Fe and Co have consecutive atomic numbers of 26 and 27, their alloys can be modeled using a virtual atom with a fractional atomic number determined by the concentration of the elements. Our goal is to study how the MAE of the system changes with Co concentration and identify promising compositions with elevated MAE. In our attempt we are motivated by previous studies, such as for FeCo, (Fe,Co)2B and (Fe,Co)16C alloys [86, 87, 88], which showed higher MAE values for intermediate compositions than for compounds at the extremes of the Fe/Co range. Moreover, we will determine how the Curie temperature and the spin and orbital magnetic moments change with Co concentration.
@@
At first, we analyze the concentration dependence of MAE of the hexa-(Fe,Co)3C phase, see Fig. 4(a) and Table 3. We observe both positive and negative MAE values, indicating uniaxial and in-plane magnetocrystalline anisotropies, respectively. The highest MAE of 1.07 MJ m-3 was obtained for a concentration of 20% Co, and the minimum MAE of -1.46 MJ m-3 for 70% Co. This trend is not surprising, as an analogous course of MAE() with a positive maximum and a negative minimum was observed for the mentioned before (Fe1-xCox)2B and (Fe1-xCox)16C alloys [87, 88]. As argued in the indicated works on uniaxial alloys, the course of the MAE() relationship is due to the filling of the valence band of the alloy as the Co concentration increases. Since MAE is a subtle quantity that depends on the details of the band structure in the close vicinity of the Fermi level, an increase in valence band filling by even about 0.1 electron per atom can lead to significant changes in MAE.
@@
For ortho-(Fe,Co)3C alloys, the Co concentration range is divided into regions: from 0% to 35%, from 35% to 75%, and from 75% to 100%, with the directions of magnetic easy axes [010], [001], and [100], respectively, see Fig. 4(b). The highest MAE value of 0.82 MJ m-3 was obtained for a concentration of 90% Co. The course of MAE() clearly differs from the result for the hexagonal phase (Fe,Co)3C and the mentioned tetragonal phases Fe-Co-B. This is due, among other things, to the presence of three main crystallographic axes, where each of them can be an easy axis with a positive MAE value. On the other hand, the details of the MAE() waveform itself are due, as we mentioned above, to the gradual filling of the valence band as the Co concentration increases. Additionally, for orthorhombic alloys, we show the energy difference between two higher energies (DE) complementing the MAE. The low DE32 values, as observed for Co-rich phases, strengthen the uniaxial character of the magnetocrystalline anisotropy of these systems.
Unfortunately, there is no systematic study of the course of the magnetocrystalline anisotropy constant on Co concentration in (Fe,Co)3C alloys to verify our theoretical predictions. From scattered information, we learn that Choe et al. obtained for the powder ortho-Fe3C sample an MAE equal to 0.405 MJ m-3 at 5 K [20]. Wu et al. synthesized a melt-spun sample of Fe2CoC and measured the MAE for it at 10 K as 0.96 MJ m-3 [17]. However, their analysis showed the presence of about 10% by weight of the bcc Fe phase in the Fe2CoC sample [17], which can affect the anisotropy constant. In contrast, our calculations for about 1/3 Co concentration indicate MAE close to zero.
A feature of our MAE calculations is that they are made for a perfect monocrystal at 0 K, and modeling chemical disorder with VCA leads to overestimation of MAE. However, the trends thus predicted turned out to be in line with experimental results [87]. The anisotropy constant of real samples can be significantly affected by numerous additional factors. In addition to the role of temperature, in the case of (Fe,Co)3C alloys we are faced with the chemical instability of the phase, the presence of microstructure and extraneous phases in the samples, and in the case of nanoparticles, the influence of nanostructuring. Given the complexity of the problem, the calculated MAEs for monocrystals at 0 K are a good starting point for understanding the relationships governing doping, and the MAE values obtained from the calculations are promising from an application point of view and encourages further exploration of magnetically hard materials in the cementite family.
@@
The total spin magnetic moment () is highest for the terminal Fe3C phases for both hexagonal and orthorhombic structures, and decrease with Co concentration, see Figs. 4(c,d). We do not observe a pronounced maximum in the intermediate range. This behavior is similar to that observed, for example, for Fe/Co monoborides [89, 68]. The observed trend has to do with an increase in the number of electrons per atom with an increase in the concentration of Co in the system. Additional electrons from Co atoms fill the valence band, especially the 3 orbitals, see Fig. 2. Since the majority Fe 3 spin channels are almost completely occupied, additional electrons fill the minority 3 spin channels, which reduces the occupancy difference between the spin channels, which is what defines the spin magnetic moment.
@@
Hexagonal (Fe,Co)3C alloys with Co concentration from 0 to 30% are magnetically semi-hard, see Fig. 4(g). For negative values of MAE, seen in the center of the concentration range, we assumed = 0, see Table 3. The hexagonal carbide with 90% Co indicates = 1.13, which defines it as magnetically hard. For orthorhombic alloys, see Fig. 4(h), semi-hard magnetic properties are observed for most of Co concentrations.
In the previous section, we discussed the Curie temperature results for the stoichiometric compounds. The calculated mean-field Curie temperatures were 1040 K for hexa-Fe3C, 310 K for hexa-Co3C, 570 K for ortho-Fe3C, and 458 K for ortho-Co3C. In Fig. 5, we show the courses of calculated as a function of Co concentration. For hexa-(Fe,Co)3C phase, the relation is close to linear, while for the ortho-(Fe,Co)3C phase a maximum around 20% Co is visible. The close to linear relationship resembles the analogous experimental results for (Fe,Co)B and (Fe,Co)2B alloys [89, 91], and the maximum in resembles results for bcc FeCo alloys [92]. As in the discussed trends for MAE and magnetic moment, the trends for depend primarily on the valence band filling of the systems under study increasing with Co concentration, see also Ref. [93]. As experiments for orhto-Co3C nanoparticles have shown [24, 25, 26, 27], we expect that for ortho-(Fe,Co)3C nanoparticles as well, Curie temperatures should be able to be raised to values above 500 K.
In summary, we calculated the basic magnetic properties for hexa- and ortho-(Fe,Co)3C alloys. Hexagonal alloys show a number of promising magnetic properties, such as a high Curie temperature of 1040 K for hexa-Fe3C, and a magnetic hardness () of 0.83 for hexa-(Fe0.8Co0.2)3C. Nevertheless, their experimental synthesis has so far presented great difficulties, as explained by the positive formation energies of hexa-Fe3C and hexa-Co3C compounds, above those for corresponding orthorhombic phases. On the other hand, it has been experimentally proven that the ortho-(Fe,Co)3C phase is stabilizable. DFT calculations predict for it decrease of spin magnetic moment with Co concentration, and maximum in Curie temperature at about 20% Co. Experimental Curie temperatures determined for orthorhombic (Fe,Co)3C nanoparticles of more than 500 K offer hope for the practical application of these alloys. Experiments also showed Curie temperature dispersion for ortho-Co3C nanoparticles, suggesting a strong effect of structuring on the magnetic properties of the system, including magnetic moment, magnetocrystalline anisotropy and Curie temperature. Most probably due to this feature, the MAE calculated here for ortho-(Fe,Co)3C alloys differ significantly from experimental data. However, at the same time, calculations and experiments indicate for selected ortho-(Fe,Co)3C compositions the promising MAE values of about 1.0 MJ m-3 at low temperatures, indicating the possibility of optimizing the alloys to reach similar values above room temperature. Mentioned optimization would consist of selecting proper Fe/Co ratio, nanoparticle size, and additions of other transition metal elements. The presence of Co in the alloy is, however, not advantageous from an economic point of view, so alloys with a high Co content are unlikely for applications requiring large amounts of material.
@@ III.3.1 Formation energy
The plots of dependencies of formation energies on 3 and 4 transition metal substitutions in ortho-Fe3C are almost identical to the previously mentioned theoretical results [37, 54]. Although the formation energies of the alloys studied provide valuable information, the idealized single-phase nature of the models means that the actual impact of additives is best verified experimentally. Bulk ortho-Fe3C has been tested by Umemoto et al. [31], who found that Cr, Mn, V and Mo additions stabilize the cementite, whereas Ti, Ni, and Si destabilize it. Furthermore, it has been determined that the substitution of 5 at.% Ti leads to the appearance of additional phases of TiC and -Fe [31]. The formation energies of ortho-Co3C alloys are higher than those of ortho-Fe3C, however, still relatively low.
@@ III.3.2 Curie temperature
In Fig. 7 we present the mean-field theory Curie temperatures () calculated for hexa-(Fe0.916X0.084)3C, ortho-(Fe0.916X0.084)3C, and ortho-(Co0.916X0.084)3C alloys with transition metals X. The transition metal concentration of 0.084 was chosen to reproduce the concentration of transition metal element X in Fe11X1C4 models.
In Fig. 7 we see, that the dependencies of on the atomic number of substitution have a clear sinusoidal trend and that each substitution reduces the of the parental hexa-Fe3C compound. Nevertheless, all the considered hexa-Fe3C + X alloys show relatively high .
Calculated trends for transition metal elements suggest that all substitutions, except Co, will significantly reduce the of the alloy. Whereas, the elements 4 and 5 will have stronger effect than the elements 3. The transition metal elements on the left side of the periods reduce to zero, which means the absence of magnetic ordering. Previous experiments confirm the reduction of Curie temperature after alloying of cementite with Cr, Mn, or Mo [29, 30]. However, it has been measured that up to about 2% Ni raises the Curie temperature of the cementite [29]. The effect of lowering can make it difficult or impossible to use most transition metals as additives to affect MAE and magnetic hardness. The solution could be, for example, a combination of transition metal alloying on the Fe site with B alloying on the C side. Indeed, the experiment showed a linear growth in up to 851 K with increasing B concentration in ortho-Fe3(B,C) borocarbides [21].
Calculations indicate that only Fe substitution raises the of the ortho-Co3C-based system, while all other elements considered lower it. The determined values below and well below 400 K suggest the impossibility of using most ortho-Co3C + X alloys as permanent magnets. The solution, as we suggested above, could be the alloying of B at the C position, since, as shown experimentally, the of ortho-Co3B is much higher than the value for ortho-Co3C and is 750 K [16].